OpenStudy (anonymous):

Write a quadratic function in standard form having zeros of -9 and -12.

OpenStudy (jack1):

y = 0 @ x = -9 and x = -12 so that gives the factors of (x+9) and (x+12) [as (-9 + 9) = 0 and (-12 + 12) = 0 so multiply out (x+9)(x+12) ricardo, that'll be ur quadratic

OpenStudy (jack1):

standard form should resemble : \[y = ax^2 + bx + c\]

OpenStudy (jack1):

u all good with that dude...?

OpenStudy (anonymous):

yeah! thanks!

OpenStudy (jack1):

sweet as bru

OpenStudy (anonymous):

So would the work look like this? \[x = -9, x = -12\] \[y = (x - 9)(x - 12)\] \[y = x^2 - 21x + 108\] @Jack1